-3b - 6 + 7b =6 what does B = ?
Answer by RAY100(1637)

-3b -6 +7b =6,,,,,combine like terms
.
4b =12
.
b=3
.
ckeck,,,-3(3) -6 +7(3) = -9-6+21 =6,,,,ok